OpenStudy (anonymous):

larry wants to buy some carpeting for his living room. the length of the room is 2 times the width and the total area of the room is 8 square meters.

OpenStudy (anonymous):

what is the legth of the living room? sorry i forgot that part.

OpenStudy (unklerhaukus):

\[l=2w;\qquad A=lw=8\]

OpenStudy (anonymous):

let the length be 'l' and width be 'w' l=2w -----(1) area of the room = l*w 8 = l*w 8 = 2w*w [from (1)] \[4 = w^2\] \[w = \sqrt{4} =2\] \[\therefore\ length\ = 2*2 = 4\ \ [\ l=2w]\]

OpenStudy (unklerhaukus):

\[w=\sqrt{4}=\pm2\] we can reject the negative answer because the width of a carpet must be positive

OpenStudy (anonymous):

thank you guys